find "x"
a(x-3)+8=b(x-1)

Answer by stanbon(75887)

find "x"
a(x-3)+8=b(x-1)
ax-3a+8 = bx-b
Get the x terms together on one side:
ax-bx = 3a-b-8
x(a-b) = 3a-b-8
x = [3a-b-8]/[a-b]
============
Cheers,
Stan H.